OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (hoblos):

x - y = - 5

OpenStudy (anonymous):

Find the slope first Slope = Rise/ Run Slope = y1-y2 / x1-x2 (-2,3) - > x1 = -2, y1 = 3 (2,7) -> x2 = 2, y2 = 7 Slope = (3-7)/(-2-2) Slope = -4 / -4 Slope = 1 Then find the y intercept y = mx + b I will use (2,7) 7 = 1(2) + b b = 7- 2 = 5 Therefore, y = x + 5 y - x = 5 or x - y = -5 (This is the answer )

OpenStudy (anonymous):

Thank you soo much ♥

OpenStudy (anonymous):

yw .. ^.^

OpenStudy (anonymous):

yeah x-y=-5 only

OpenStudy (hoblos):

well here it was much easier than that you only had to replace (-2,3) in the equations and you will notice that it is true only for x-y=-5